# PriceFM: Foundation Model for Probabilistic Electricity Price Forecasting

### v1

@@ Introduction
For the upper quantile ($\tau=0.9$), a residual price trajectory $\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}$ is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer:
@@
where a non-negative function $g(\cdot)$, such as absolute-value function, is applied to the price residual. The upper quantile forecast is obtained by adding this non-negative residual to the median prediction:
@@
This hierarchical design guarantees that the upper quantile prediction is greater than or equal to the lower one at each time step, effectively overcoming quantile crossing.
Loss
We introduce the Average Quantile Loss (AQL) as the training objective for multi-region, multi-timestep, and multi-quantile probabilistic forecasting. Let $y_{i,{r_{\mathrm{out}}},t}$ denote the ground-truth price for the $i$-th training sample, output region $r_{\mathrm{out}}$}, and timestep $t$, and let $\hat{y}_{i,{r_{\mathrm{out}}},t,\tau}$ be the corresponding predicted quantile. The AQL across all training samples, regions, timesteps, and quantiles is computed as:
@@
where $N$ is the number of samples, and the quantile loss function $L_{\tau}$ for a single prediction task is defined as:
@@
A lower AQL indicates more accurate probabilistic forecasting across spatial and temporal dimensions.
Other Details
We use the Adam optimizer with an initial learning rate of $4\times 10^{-3}$, which decays exponentially by a factor of 0.95 every 10 epochs. The batch size is 8, which introduces a slight regularization effect. The model is trained for 50 epochs, and the checkpoint with the lowest validation loss is saved.
Experiment
@@
For the three-year dataset, we use the first two years of the dataset (67%) for training, and the final year (33%) is evenly split into validation and testing sets, strictly preserving the temporal order to ensure realistic out-of-sample evaluation. We assess the model performance using the quantile loss at individual levels (Q10, Q50, and Q90), AQL, Average Quantile Crossing Rate (AQCR), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and coefficient of determination ($R^{2}$). The Diebold-Mariano (DM) test is applied to determine if two models have a significant difference in the probabilistic and pointwise forecasting (Diebold and Mariano 2002). Each metric is explained in Appendix, Section Metrics.
Model Comparison
We include several spatial and temporal models as baselines. Model descriptions and hyperparameter optimization are detailed in Appendix, Section Baselines.
Spatial Models
Following (Yang et al. 2024; Meng et al. 2024), where variants of GNNs, such as the Graph Convolutional Network (GraphConv) and the Graph Attention Network (GraphAttn), have made successes in electricity price forecasting, we include them as baselines. We also compare with other advanced variants, such as a GNN based on sample-and-aggregate operations (GraphSAGE) (Hamilton, Ying, and Leskovec 2017), a diffusion-based GNN (GraphDiff) (Li et al. 2018), and a GNN with Auto-Regressive Moving Average filters (GraphARMA) (Bianchi et al. 2021).
@@
Results in Table 1 show that PriceFM statistically outperforms all spatial models, confirmed by both the probabilistic and pointwise DM test. A common limitation of these models is that they have only limited spatial priors, relying primarily on data-driven learning to assign spatial importance. Without an explicit mechanism to enforce spatial decay, they lack regularization over noisy regions. As a result, they require large datasets to generalize. However, despite our dataset spanning three years, which is considered large in the domain, the daily forecasting requirement limits the training set to only around 700 samples, making it unsuitable for such models.
Model Q10 $\downarrow$ Q50 $\downarrow$ Q90 $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ ${R}^{2}$ $\uparrow$
GraphConv 4.81±0.16 10.76±0.16 7.11±0.33 7.56±0.15 2.19±0.76 35.75±0.60 21.53±0.33 0.51±0.02
GraphAttn 5.08±0.31 11.58±0.37 7.90±0.40 8.18±0.29 1.41±0.47 37.63±0.62 23.15±0.75 0.44±0.01
@@
TimeXer 47.38±1.18 28.29±0.71 0.14±0.02
PriceFM 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
Table 1: Model performance comparison on the testing set. The symbol “–” indicates that the model does not support probabilistic forecasting by design. All metrics are reported as mean±standard deviation over 5 independent runs. The best results are shown in bold, and the second-best are underlined. The units of Q10, Q50, Q90, AQL, RMSE, and MAE are expressed in €/MWh (denormalized), and AQCR in %.
Temporal Models
We include FEDFormer (Zhou et al. 2022), TimesNet (Wu et al. 2023), and TimeXer (Wang et al. 2024) in our evaluation. As these foundation models do not support graph-based input, features from all regions are concatenated along the feature dimension. Including these models aims to investigate whether these pure time-series models can capture spatial patterns without prior graph knowledge.
Results in Table 1 show that PriceFM statistically outperforms all temporal models, confirmed by the pointwise DM test. In particular, the inclusion of noisy features from weakly related regions easily leads to overfitting, given the high model complexity of these time-series foundation models. This observation confirms that the temporal models struggle to recognize spatial patterns without explicit injection of graph knowledge. In contrast, PriceFM incorporates a graph decay mechanism that acts as a spatial regularizer, thereby attenuating the influence of noisy regions.
Ablation Study
We systematically evaluate the impact of key hyperparameters and architectural design choices, including spatiotemporal configurations, fusion strategies in the fusion block, graph decay mechanisms in the graph block, and variants of the hierarchical quantile head.
Spatiotemporal Configurations
We assess the impact of two key hyperparameters that control the spatial decay strength and temporal input length:
  • Curvature Parameter: Spatially, we evaluate $c\in\{-1.0,-0.8,\ldots,0.8,1.0\}$ in increments of 0.2, ranging from weak decay to strong decay. In total, 2,090 trials are conducted to determine the optimal curvature value for each output region individually, based on validation loss.
@@
Figure 5 illustrates the testing loss and the distribution of optimal curvature values across all regions. Most regions confirm spatial interdependencies ($c\neq 1.0$). The identified optimal curvature shows a weak negative correlation ($\rho=-0.18$) with the number of directly connected regions, mentioned in Figure 1. The results in Table 2 indicate that the optimal backward-looking window size is 24, potentially because information from the distant past becomes outdated.
Figure 5: Spatial distribution of testing loss and curvature values. a Average quantile loss per region on the testing set. Western European regions exhibit lower losses, whereas BG, HU, and RO show particularly high losses (orange areas). b Optimal curvature value per region. Particularly, 8 regions from IT and SE have a curvature value of 1, indicating optimal performance by excluding neighboring features.
Method Q10 $\downarrow$ Q50 $\downarrow$ Q90 $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ ${R}^{2}$ $\uparrow$
$L=24^{\dagger}$ 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
@@
Table 2: Ablation studies of different module choices: temporal configuration (rows 1-3), fusion block (rows 4–6), graph block (rows 7–9), and hierarchical quantile head (rows 10–13). The symbol marks the method used in PriceFM. All metrics are reported as mean±standard deviation over 5 independent runs. The best results are shown in bold, and the second-best are underlined. The units of Q10, Q50, Q90, AQL, RMSE, and MAE are expressed in €/MWh (denormalized), and AQCR in %.
Fusion Block
We compare the residual addition from the fusion block, described in Eq. (4), with two fusion strategies:
  • Concatenation: We replace Eq. (4) by first flattening both $\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, then concatenating them:
@@
The results in Table 2 show that replacing the residual addition with concatenation leads to 6.10% higher AQL for probabilistic prediction. Switching to cross-attention yields comparable performance to residual addition, while introducing additional model parameters and unnecessary complexity. This suggests that the residual addition strikes a favorable balance between performance and model simplicity.
Graph Block
We compare the graph decay mask from the graph block, described in Eq. (8), with two alternatives:
  • Random Graph Decay Mask: We replace Eq. (8) with a randomly sampled vector, where each decay weight is drawn independently from a uniform distribution over $[0,1]$, thereby removing the spatial graph prior:
@@
  • No Graph Decay: We remove the decay mask, which simplifies Eq. (9) to a uniform average over input regions:
@@
The results in Table 2 demonstrate that randomizing or removing the graph decay mask, which discards the prior graph knowledge, leads to a significant drop in all metrics. We also observe that such results, obtained without prior graph knowledge, are on par with those of GNN variants, shown in Table 1. We emphasize that relying on pure data-driven learning without an explicit decay mechanism leads to a loss of the key inductive bias, limiting the model’s performance, especially when the training data is scarce.
Hierarchical Quantile Head
We evaluate the impact of different non-negative functions used in Eq. (12) and (14), and include a standard multi-quantile head for comparison.
  • Non-Negative Functions: We replace the absolute-value function with either a square function or ReLU:
@@
  • Standard Multi-Quantile Head: The Eq. (11) and (13) are skipped, and the shared representation $\mathbf{U}_{r_{\mathrm{out}}}$ is passed directly to three independent dense layers to produce each quantile trajectory:
    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.1}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (23)
    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.5}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (24)
    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.9}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}}.$ (25)
@@
Conclusion
In this paper, we introduced and released a dataset, which will benefit both the research community and the energy industry. We proposed and released PriceFM, a foundation model tailored for interconnected European electricity markets. Extensive experiments and ablation studies demonstrate that PriceFM outperforms competitive baselines and highlight the importance of spatial context. By enabling more accurate and comprehensive probabilistic electricity price forecasting, our work has the potential to support better decision-making in energy trading and grid management, contributing to a more efficient and resilient energy ecosystem.
Despite these promising results, our study has two key limitations. First, the graph decay function is empirically designed; the performance may improve with alternative functions. Second, PriceFM relies on the current transmission topology, which may become outdated as transmission networks evolve, requiring retraining to adapt to new scenarios.
Appendix
Code Guideline
@@ Conclusion
Step 1: Create a folder named PriceFM, along with subfolders Data, Model, and Result. Place the energy data EU_Spatiotemporal_Energy_Data.csv into Data, and place PriceFM.py inside the PriceFM folder.
@@
Step 3: After execution, check the Model and Result folders for the saved model weights and evaluation metrics.
Hardware and Computation
The PriceFM is evaluated on both an NVIDIA A100 GPU and an Intel Core i7-1265U CPU, respectively. The NVIDIA A100 is designed for high-performance computing and deep learning workloads, offering 80 GB of high-bandwidth memory and up to 6,912 CUDA cores. In contrast, the Intel i7-1265U is a power-efficient CPU commonly found in standard laptops. This dual evaluation setup allows us to assess the feasibility of deploying our model in environments without access to advanced GPU hardware.
Under the training setup described in Section Model, the training time is approximately 4–5 minutes on the A100 GPU and 12–13 minutes on the i7 CPU. Due to the small batch size used, the GPU does not achieve a significant speedup over the CPU. Inference time for both setups is under 10 seconds. We note that neither training nor inference time is critical for our application, as bid submissions can occur at any point before the market gate closure.
Lookup Table and Feature Availability

### v2

@@ 4.3 Head
For the upper quantile ($\tau=0.9$), a residual price trajectory $\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}$ is produced from $\mathbf{U}_{r_{\mathrm{out}}}$:
@@
where a non-negative function $g(\cdot)$, such as absolute-value function, is applied to the price residual. The upper quantile forecast is obtained by adding this non-negative residual to the median:
@@
This hierarchical design guarantees that the upper quantile prediction is greater than or equal to the lower one at each time step, overcoming quantile crossing.
4.4 Loss
We introduce the Average Quantile Loss (AQL) as the training objective for multi-region, multi-timestep, and multi-quantile probabilistic forecasting. Let $y_{i,{r_{\mathrm{out}}},t}$ denote the ground-truth price for the $i$-th training sample, output region $r_{\mathrm{out}}$}, and timestep $t$, and let $\hat{y}_{i,{r_{\mathrm{out}}},t,\tau}$ be the corresponding predicted quantile. The AQL is computed as:
@@ 4.4 Loss
where $N$ is the number of samples, and the quantile loss $L_{\tau}$ is defined as:
@@
5 Experiment
@@
We split the data into training (2022-01-01 to 2024-01-01), validation (2024-01-01 to 2024-07-01), and testing (2024-07-01 to 2025-01-01). The choice of the testing period aims to include numerous extreme prices, as illustrated in Figure 2 (a). We assess the model performance using the quantile losses ($\text{Q}_{0.1}$, $\text{Q}_{0.5}$, and $\text{Q}_{0.9}$), AQL, Average Quantile Crossing Rate (AQCR), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Coefficient of Determination ($\text{R}^{2}$). The Diebold-Mariano (DM) test is applied to determine if two models have a significant difference Diebold & Mariano (2002). All metrics are explained in Appendix G. The hyperparameters are detailed in Appendix D
5.1 Model Comparison
5.1.1 Naïve Baselines
We include three seasonal naïve baselines as reference models, where only historical prices are used as input: Naïve1 uses 24 prices from the previous day; Naïve2 uses 24 prices averaged over the past three days; Naïve3 uses 24 prices averaged over the past seven days. To obtain probabilistic results, we compute empirical quantiles at individual levels ($\mathcal{Q}=\{0.1,0.5,0.9\}$) for each delivery hour. The seasonal naïves are commonly used to evaluate the autoregressive strength of the signal and often serve as strong baselines Ziel & Weron (2018); Lago et al. (2021).
The results from Table 1 show that PriceFM significantly outperforms the naïve baselines, confirmed by both the probabilistic and pointwise DM tests, with all $p\text{-values}<0.05$ and negative DM values. Specifically, the AQL values of baselines are between 36.83% and 44.80% higher, with an AQCR of 0.00%, as these forecasts are directly computed from historical values. Moreover, the high RMSE and MAE, together with the low $\text{R}^{2}$, observed in the naïve baselines suggest limited performance.
5.1.2 Graph Models
We compare with multiple GNN variants: Graph Convolutional Network (GCN) Kipf (2016), Graph Attention Network (GAT) Veličković et al. (2017), GraphSAGE Hamilton et al. (2017), GraphDiffusion Li et al. (2018), and GraphARMA Bianchi et al. (2021). The adjacency matrix of these graph models is explained in Appendix E.
@@ 5.1.2 Graph Models
From Table 1, we observe that PriceFM outperforms all graph models, confirmed by both the probabilistic and pointwise DM tests, with all $p\text{-values}<0.05$ and negative DM values. Notably, PriceFM reduces the AQL by between 8.86% and 15.77% and consistently achieves 0.00% AQCR compared to graph baselines. A common limitation of these graph baselines is that they lack explicit regularization over noisy regions and rely primarily on data-driven learning to assign spatial importance. As a result, they require large datasets to generalize. Despite our dataset spanning three years, which is considered large in the domain, the daily forecasting requirement limits the training set to only around 700 samples, making it unsuitable for such models.
Table 1: Model comparison on the testing set. The symbol “–” indicates that the model does not support probabilistic forecasting by design. The naïve baselines have zero standard deviation, as they are deterministically computed. All metrics are reported as mean±standard deviation over 5 independent runs. The best results are shown in bold, and the second-best are underlined. The units of $\text{Q}_{0.1}$, $\text{Q}_{0.5}$, $\text{Q}_{0.9}$, AQL, RMSE, and MAE are expressed in €/MWh, and AQCR in %.
Model $\text{Q}_{\textbf{{0.1}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.5}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.9}}}$ $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ $\textbf{R}^{\textbf{2}}$ $\uparrow$
Naïve1 5.68±0.00 14.30±0.00 8.31±0.00 9.43±0.00 0.00±0.00 46.90±0.00 28.60±0.00 0.16±0.00
Naïve2 6.01±0.00 14.42±0.00 8.44±0.00 9.62±0.00 0.00±0.00 46.73±0.00 28.84±0.00 0.17±0.00
Naïve3 6.65±0.00 14.58±0.00 8.71±0.00 9.98±0.00 0.00±0.00 46.73±0.00 29.15±0.00 0.17±0.00
GCN 4.81±0.16 10.76±0.16 7.11±0.33 7.56±0.15 2.19±0.76 35.75±0.60 21.53±0.33 0.51±0.02
GAT 5.08±0.31 11.58±0.37 7.90±0.40 8.18±0.29 1.41±0.47 37.63±0.62 23.15±0.75 0.44±0.01
@@
TimeXer 44.57±0.66 26.52±0.55 0.30±0.01
PriceFM 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
5.1.3 Time-Series Foundation Models
We include several time-series foundation models: FEDFormer Zhou et al. (2022), iTransformer Liu et al. (2023), PatchTST Nie et al. (2023), TimesNet Wu et al. (2023), and TimeXer Wang et al. (2024), to investigate whether these pure time-series models can capture spatial patterns without prior graph knowledge. As these foundation models do not support graph-based input, features from all regions are concatenated along the feature dimension.
The results in Table 1 show that these time-series foundation models achieve similar RMSE and MAE but exhibit better explained variance, as indicated by higher $\text{R}^{2}$, compared to the naïve baselines. Notably, PriceFM outperforms all time-series foundation models, confirmed by the pointwise DM test, with all $p\text{-values}<0.05$ and negative DM values. On average, PriceFM improves RMSE, MAE, and $\text{R}^{2}$ by 27.98%, 26.38%, and 0.31, respectively. Given the high complexity of these foundation models, the inclusion of features from noisy regions easily leads to overfitting. This observation confirms that the pure time-series models struggle to recognize useful spatial patterns. In contrast, PriceFM incorporates a graph decay mechanism that acts as a spatial regularizer, thereby attenuating the influence of noisy regions.
5.2 Ablation Study
5.2.1 Spatiotemporal Configurations
  • Curvature Parameter: Spatially, we evaluate $c\in\{-1.0,-0.8,\ldots,0.8,1.0\}$ in increments of 0.2, ranging from weak decay to strong decay. In total, 2,090 trials are conducted to determine the optimal curvature value for each output region individually.
@@ 5.2.1 Spatiotemporal Configurations
Spatially, Figure 4 illustrates the testing loss and the distribution of optimal curvature values across all regions. Most regions confirm spatial interdependencies ($c\neq 1.0$). Temporally, the results in Table 2 indicate that the optimal backward-looking window size is 24, potentially because information from the distant past becomes outdated.
Figure 4: Spatial distribution of testing loss and curvature values. (a) Average quantile loss per region on the testing set. Western European regions exhibit lower losses, whereas BG, HU, and RO show particularly high losses (orange areas). (b) Optimal curvature value per region. Notably, BG, ES, FR, FI, and partial regions from IT and SE have a curvature value of $1.0$, indicating optimal performance by excluding neighboring features.
Table 2: Ablation studies of different module choices: temporal configuration (rows 1-3), fusion block (rows 4–6), graph block (rows 7–9), and hierarchical quantile head (rows 10–13). The symbol marks the method used in PriceFM.
Method $\text{Q}_{\textbf{{0.1}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.5}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.9}}}$ $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ $\textbf{R}^{\textbf{2}}$ $\uparrow$
$L=24^{\dagger}$ 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
@@
5.2.2 Fusion Block
  • Concatenation: We replace Equation 4 by first flattening both $\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, and then concatenating them:
@@ 5.2.2 Fusion Block
The results in Table 2 show that replacing the residual addition with concatenation leads to 6.10% higher AQL for probabilistic prediction. Switching to cross-attention yields comparable performance to residual addition, while introducing additional model parameters. This suggests that the residual addition strikes a favorable balance between predictive performance and model simplicity.
5.2.3 Graph Block
  • Random Graph Decay Mask: We replace Equation 8 with a randomly sampled vector, where each decay weight is drawn independently from a uniform distribution over $[0,1]$, thereby removing the spatial graph prior:
@@ 5.2.3 Graph Block
  • No Graph Decay: We remove the decay mask, which simplifies Equation 9 to a uniform average over input regions:
@@
The results in Table 2 demonstrate that randomizing or removing the graph decay mask, which discards the prior graph knowledge, leads to a significant drop in all metrics. We also observe that such results are on par with those of GNN baselines. We emphasize that relying on pure data-driven learning without an explicit decay mechanism leads to a loss of the key inductive bias, limiting the model’s performance, especially when the training data is scarce.
5.2.4 Hierarchical Quantile Head
  • Non-Negative Functions: We replace the absolute-value function used in Equation 12 and 14 with either a square function or ReLU:
@@ 5.2.4 Hierarchical Quantile Head
  • Standard Multi-Quantile Head: The Equation 11 and 13 are skipped, and $\mathbf{U}_{r_{\mathrm{out}}}$ is passed directly to three independent dense layers to produce quantile trajectories.
@@
6 Conclusion
In this paper, we introduced and released a dataset, which will benefit both the research community and the energy industry. We proposed and released PriceFM, a foundation model with a generic architecture applicable to all European electricity markets without relying on pretraining, similar to time-series foundation models used in this work. Extensive experiments and ablation studies demonstrate that PriceFM outperforms competitive baselines and highlight the importance of spatial context. By enabling more accurate and comprehensive probabilistic electricity price forecasting, our work has the potential to support better decision-making in energy trading and grid management.
Several directions remain for future work. First, as the current graph decay function is empirically defined, exploring alternative formulations could improve spatial representations. Second, as the transmission network evolves, model retraining may be required to account for structural changes. Third, since our dataset is sampled at an hourly resolution, we aim to extend PriceFM to support quarter-hourly forecasts, in anticipation of more European bidding zones transitioning to 15-minute markets in the coming years.
7 Reproducibility Statement
To support reproducibility, we release well-documented code with an easy-to-use three-step pipeline, along with the code structure and usage guidelines in Appendix A. We report hardware and runtime details in Appendix B to facilitate realistic deployment of our proposed model. As the release and cleaning of the dataset are part of our contributions, the details are described in Section Data and Appendix C. Additional reproducibility details are also provided, including data scaling (Appendix F), evaluation metrics (Appendix G), hyperparameters and training procedures (Appendix D).
Appendix A Code Guideline
@@
Step 1: Create a folder named PriceFM, along with subfolders Data, Figure, Model, and Result. Place the energy data EU_Spatiotemporal_Energy_Data.csv into Data, and place PriceFM.py inside the PriceFM folder.
@@
Step 3: After execution, you can inspect: Figure/ for visualizations of forecasts versus true prices; Model/ for saved model weights; Result/ for evaluation metrics.
Optional: To better understand the code structure and functionality, run Tutorial.ipynb block by block.
Appendix B Hardware and Computation
The PriceFM is evaluated on both an NVIDIA A100 GPU and an Intel Core i7-1265U CPU, respectively. The NVIDIA A100 is designed for high-performance computing and deep learning workloads, offering 80 GB of high-bandwidth memory and up to 6,912 CUDA cores. In contrast, the Intel i7-1265U is a power-efficient CPU commonly found in standard laptops. Under the training setup described in Section Model, the training time is approximately 4–5 minutes on the A100 GPU and 12–13 minutes on the i7 CPU. Inference time for both setups is under 10 seconds. We note that neither training nor inference time is critical for our application, as bid submissions can occur at any point before the market gate closure on a daily basis.
Appendix C Lookup Table and Feature Availability